DIFFERENTIAL GEOMETRY/PDE SEMINAR

Wednesday, May 24 Padelford C-36 3:45 pm

Hamiltonian group actions

Yael Karshon (The Hebrew University of Jerusalem, visiting UC Berkeley)

Hamiltonian group actions arise in many different contexts: classical mechanics, representations of Lie groups, and even combinatorics. We will discuss classification results for torus actions when the ”complexity” is low. The ”complexity” measures how far the action is from being ”completely integrable. This study is a joint project with Sue Tolman.